OpenStudy (anonymous):

A bird sitting 16 feet above the ground in an apple tree dislodges an apple. After how many seconds does the apple land on the ground? (Assume no branches interfere with its fall.)

OpenStudy (gudden):

 Do you know The first equation of motion ?

OpenStudy (gudden):

@Chrisgoblin there?

OpenStudy (anonymous):

yea and no

OpenStudy (gudden):

Anyway make sure that 1 foot = 0.3 m

OpenStudy (gudden):

Well.. we got to find time => time t = ? Distance/ displacement has been given = s = 16 foot = 16*.3 = 4.8m And acceleration due to gravity = 9.8 m/s^2 = a We do not have the final velocity ' v ' So, we would be using second equation of motion which is \[s = ut + \frac{ 1 }{ 2 }at ^{2}\]

OpenStudy (gudden):

And note that when an object starts to fall the initial velocity i.e. u is taken to be 0

OpenStudy (gudden):

So, plug in the values of all the known elements

OpenStudy (anonymous):

im lost...

OpenStudy (anonymous):

I am trying to base this off what my book is saying and this is what I got -4t^2+vt+s

OpenStudy (gudden):

What does your book say ?? So, that I can explain you the way it wants you to understand

OpenStudy (anonymous):

The books example doesn't explain much - A window washer drops a wet sponge from a height of 64 feet. After how many seconds does the sponge land on the ground? Use the vertical motion model to write an equation for the height h (in feet) of the sponge as a function of the time t (in seconds) after it is dropped

OpenStudy (gudden):

Alright, so you are supposed to solve it using equations of motion :)

OpenStudy (anonymous):

h=-16t^2+vt+s 0=-16t^2 +(0)t+64 0=-16(t^2-4) 0=-16(t-2)(t+2) t-2=0 or t+2=0 t=2 or t=-2 The sponge lands on the ground 2 seconds after it dropped.

OpenStudy (gudden):

Ok..!

OpenStudy (anonymous):

how did the book come up with that answer?

mathslover (mathslover):

\(t^2 -4\) = (t-2)(t+2) The book used (a-b) ( a + b) = \(a^2 - b^2\) identity.

OpenStudy (gudden):

The vertical motion formula is used in many mathematical problems. The formula is h=-16t^+vt+s. H is the final height, T is time, V is velocity, and S is starting height. (ask.com)

mathslover (mathslover):

And so, -16(t-2)(t+2) = 0 16 is constant, so, it can not be zero. t - 2 and t + 2 can be zero. for t-2 = 0 t = 2 for t + 2 = 0 t = - 2 t can never be negative so, t = 2

OpenStudy (gudden):

Why don't you, simply plug in the values in the formula h=-16t^2+vt+s. ?

OpenStudy (anonymous):

because I don't know how to do that

OpenStudy (gudden):

Alright! So listen h = Final Height = It's on the ground now and came there after traveling 16 foot.... so what can it be ? Don't we have that ? S = Starting height.. what was it ? V = velocity with which the object starts it's motion....And as I already told you that an object falling downwards is said to have a velocity of 0m/s when it starts ,...so V = 0

OpenStudy (gudden):

Can you now locate the values of these .. ? And don't worry, I'm here if any problem comes..:)

OpenStudy (anonymous):

0=-16t^2+(0)+16 ?

OpenStudy (gudden):

You figured that out !! Already! And that's correct.. :)

OpenStudy (anonymous):

thanks for the help

OpenStudy (gudden):

You are welcome..! And just cos of your question, I learnt a new concept...:)

OpenStudy (gudden):

Thanks to you too :)